OpenStudy (lilday12):

Choose the ratio that you would use to convert 880 miles to yards. Remember that there are 1,760 yards in one mile.

OpenStudy (instagrammodel):

If you want to convert 2.4 hours into minutes, you would multiply 60 minutes/1hour to it and get 144 minutes.

OpenStudy (instagrammodel):

I'm not sure if i'm 100% Right! :/

OpenStudy (mathmale):

\[miles~ \rightarrow~yards\]

OpenStudy (mathmale):

Remember that 1,760 yards in one mile, so the equivalent conversion factors derived from this identiy are \[\frac{ 1760~yards }{ 1~mile }=1~and~\frac{ 1~mile }{ 1760 yards }=1\]

OpenStudy (instagrammodel):

Ohh.

OpenStudy (mathmale):

We want to get rid of the unit "miles" in favor of "yards." Example: 2 miles = ? yards Multiply 2 miles by the conversion factor \[\frac{ 1760~yards }{ 1~mile }=1\] We know this is the correct factor to use because "miles" will cancel and leave us with "yards." \[\frac{ 2~miles }{ 1}*\frac{ 1760~yards }{ 1~mile }=3520~yards.\]

OpenStudy (mathmale):

Now, would you plese follow this example, and convert 880 miles to yards.

OpenStudy (mathmale):

@lilday12: Would you kindly inform those who are helping you when you decide to get off OpenStudy right in the middle of a conversation. I get the impression that you're not interested or don't have the patience to follow through on your own post.

OpenStudy (instagrammodel):

So, You don't wish to have more help?

OpenStudy (lilday12):

I'm sorry and thx u help me understand it kinda

OpenStudy (lilday12):

Choose the ratio that you would use to convert 2.5 centimeters to meters. Remember that there are 100 centimeters in one meter.